# Stimulating the Diffusion Model for Image Denoising via Adaptive Embedding and Ensembling

### v3

@@ 1 Introduction
As smartphones have become ubiquitous, the pursuit of capturing high-quality images has become notably more demanding. Yet, when capturing images under challenging conditions, such as low-light environments, substantial information would be lost due to imaging noise. Consequently, the field of image denoising, with a focus on achieving low distortion and high perception, has been as a vital and thriving area of research. Traditional methods [1, 2, 3] rely on image priors to guide the denoising process, but their effectiveness is limited under extreme conditions. With the development of deep learning, discriminative methods have become the mainstream method, which are usually trained by pixel-level losses [4, 5, 6, 7, 8]. Actually, the pixel-level losses tend to predict the median (or average) of all possible values rather than the realistic images [9, 10]. Thus, these discriminative methods often struggle with perceptual quality, particularly under extreme noise levels. To improve perceptual quality, other image restoration tasks usually employ generative methods. Current state-of-the-art solutions usually rely on Generative Adversarial Networks (GANs) [11, 10, 12]. However, little work has been done to improve the perceptual quality of image denoising, even though generative approaches have been introduced to address real noise modeling [13, 14]. Moreover, GAN-based methods are widely blamed for artifacts and inconsistency, leading to significant distortion, especially for extreme degradation.
Figure 1: From the denoising perspective, the reverse process of a diffusion model can be viewed as a coarse-to-fine iterative denoising process.
Recently, the diffusion model [15, 16, 17] has achieved state-of-the-art (SOTA) performance in various tasks. The methods that utilize diffusion models exhibit higher perceptual quality and fewer artifacts. Therefore, we believe that the diffusion model holds significant potential for image denoising. However, stimulating the diffusion model for image denoising still remains several critical problems. Specifically, the diffusion model is designed to receive pure Gaussian noise as input, which is different from the noisy image input required for image denoising. Furthermore, the content difference between the generated image and the desired denoised image introduces distortion. The input inconsistency problem and the content inconsistency problem are crucial to be tackled to stimulate the diffusion model for image denoising.
@@
In this paper, we propose a novel strategy to stimulate the Diffusion Model for Image Denoising (DMID), consisting of an adaptive embedding method and an adaptive ensembling method. Our insight comes from the denoising perspective towards the diffusion model. From the denoising perspective, the diffusion model and the iterative denoising methods share the same structure and similar functionalities. As illustrated in Figure 1, the diffusion model can be viewed as an iterative denoising method. This new perspective provides us with huge potential to employ a pre-trained unconditional diffusion model to tackle the aforementioned problems, which also eliminates the resource-intensive and time-consuming demands associated with training a diffusion model [18]. Firstly, the problem of input inconsistency is essentially an embedding problem. Here the noisy image and iterative denoising process correspond to the intermediate state and the reverse process subsequence of the pre-trained unconditional diffusion model, respectively. Therefore, we propose an adaptive embedding method that embeds the noisy image into an intermediate state of the pre-trained unconditional diffusion model. Our embedding method enables to perform Gaussian and real-world image denoising. Secondly, the problem of content inconsistency fundamentally constitutes a stochasticity constraint problem. Since the distortion of the diffusion model mainly comes from the stochasticity inherent in the iterative process. To tackle this problem, we propose an adaptive ensembling method that can adjust distortion and perception without additional training. Our ensembling method enables to traverse through the perception-distortion curve [19] as shown in Figure 2, which achieves SOTA performance on both distortion-based and perception-based metrics. In summary, our contributions are as follows:
@@
The organization of this paper is as follows. Firstly, we provide a brief introduction to the development of image denoising and the diffusion model in Section 2. Next, we offer a rethinking and understanding of the diffusion model from the denoising perspective in Section 3. Then, we propose our DMID strategy to tackle the existing problems in Section 4. After that, we demonstrate the performance and evaluate the effectiveness of our strategy in Section 5. Finally, we conclude this paper and point out the future directions in Section 6.
@@ 2.1 Image Denoising
Image denoising is an essential task in computer vision. It aims at restoring a clean image from its noisy counterpart. However, achieving high perception quality and low distortion in image denoising remains a challenging task. To tackle this problem, image denoising methods can be broadly divided into two categories: distortion-based methods and perception-based methods. Distortion-based methods consist of traditional denoising methods and discriminative methods. Traditional image denoising methods focus on the usage of image priors such as sparsity [1, 20], low rank [2], self-similarity [21, 22], and smoothness [3, 23]. However, the ability of these priors in extreme conditions is ultimately limited, making it difficult to adequately denoise the corrupted images. With the development of deep learning, discriminative models [4, 5, 24, 25, 6, 7, 8, 26, 27], have become the mainstream methods. The main idea of discriminative methods is to use neural networks to learn mappings from noisy images to clean images through paired noisy-clean data or just noisy data [28, 29, 30, 31, 32]. $L_{1}$ loss and $L_{2}$ loss are most widely adopted during training. However, it is well known that these pixel-level losses tend to predict the median (or average) of all possible values rather than the realistic images [9, 10], resulting in over-smooth images with poor perceptual quality. In addition, some distortion-based methods [33, 34, 35, 1] employ an iterative framework to optimize denoising performance. The iterative denoising methods usually predict a rough image and refine it through multiple iterations. The iterative denoising framework serves as a regularization and constraint. In addition, the iterative denoising framework decomposes image denoising into a series of sub-problems, making denoising easier. However, these methods still struggle with perceptual quality. In contrast to distortion-based methods, perception-based methods typically exhibit superior perceptual quality but have significant distortion. Perception-based methods usually refer to generative methods. Current perception-based approaches with competitive performance for other image restoration usually rely on Generative Adversarial Networks (GAN) [11, 10, 12]. However, GAN-based methods often introduce artifacts and inconsistent details that are not present in the original clean image. The artifacts usually lead to significant distortion, especially for extreme degradation. As for image denoising, many methods employ GAN and flow [36] to solve real noise modeling problems [13, 37, 14, 38, 39, 40, 41] to get paired data easily. It is really a shame that little work has attempted to improve the perceptual quality of image denoising.
@@ 2.2 Diffusion Model
The diffusion model is a burgeoning likelihood-based generative model and has demonstrated remarkable success over other models in various tasks [18, 15, 42]. The diffusion model is originally proposed separately from diffusion-based [16] and score-matching-based [43] perspectives. Later Ho et al. [15] establish a connection between the diffusion model and denoising score matching with Langevin dynamics and demonstrates its excellent performance and enormous potential. Recently, Dhariwal et al. [18] improves unconditional image synthesis quality by finding a better architecture and firstly proposes a kind of conditional diffusion model. This work successfully brings diffusion models back to the public view. After that, conditional diffusion models adapt to various tasks and have achieved brilliant performances in a series of image restoration tasks, such as image super-resolution [44, 45], image inpainting [46], image deblurring [47]. Recently, some methods [48, 49] have exclusively employed pre-trained unconditional diffusion models, originally trained for image synthesis, to address linear image restoration problems. These methods typically assume that the observed image $y$ is degraded as $y=Hx+n$, where $H$ represents the degradation matrix, $n$ denotes noise, and $x$ signifies the desired clean image. The existing methods usually replace some content with the original degraded image and preserve the others during each sampling step of the diffusion model. These methods achieve this by decomposing the degradation matrix $H$ to identify the boundary of the reserved information area. The popular techniques are Singular Value Decomposition and Range-Null Space Decomposition.
@@ 3.1 Diffusion Model Pipeline
In this section, we give a rough introduction to the diffusion model. Diffusion models are composed of a T-timestep forward process and a T-timestep backward process. The forward process, also known as the diffusion process, gradually adds i.i.d. Gaussian noise to the initial data $x_{0}$ and converges to the unit spherical Gaussian distribution. The reverse process is the inference process, which starts with pure Gaussian noise $x_{T}\sim\mathcal{N}(0,1)$ and eventually obtains the expected high-quality image $x_{0}$. The forward process is changed from the previous state $x_{t-1}$ follows the Markov chain to generate current state $x_{t}$:
@@
where $\alpha_{t}=1-\beta_{t}$, $\bar{\alpha_{t}}=\prod_{i=1}^{t}(1-\beta_{i})$. The reverse process generates an image by a series of sampling processes from $x_{t}$ to $x_{t-1}$:
@@
where $\mu_{\theta}(x_{t},t)$ is modeled by the neural network. The reverse process gradually transforms the Gaussian distribution $x_{T}$ into the expected data distribution $x_{0}$. Different sampling strategies of eq. 3 have been proposed from DDPM [15] to DDIM [51]. These strategies can be summarized as a predicted item and an additional noise item:
@@
3.2 Iterative Denoising Framework
In this section, we point out the iterative framework for image denoising. Some distortion-based methods [33, 34, 35, 1] employ an iterative framework to optimize denoising performance. These methods usually predict a rough image and refine it through many iterations. Such an iterative framework breaks down the problem into a series of sub-problems, leading to a coarse-to-fine denoising process. The iterative denoising framework can be developed as Algorithm 1. Specifically, the inference of this framework includes several iterations. In each iteration, a rough denoising result $\hat{x}_{0}$ is firstly estimated from the current image $x_{t}$. After that, a weighted original noisy image $y$ is introduced to obtain a corrected version $x_{t-1}$ with lower noise than the previous image $x_{t}$, as:
$x_{t-1}=\hat{x}_{0}+\gamma_{t}(y-\hat{x}_{0})\ .$ (5)
@@ 3.2 Iterative Denoising Framework
2:$skip=\text{floor}({N/S_{t}})$
3:$\hat{x}=\textbf{0}$
4:for $i=1,\dotsc,R_{t}$ do
5:    for $t$ in reversed(range$(0,N,skip)$do
6:         $\epsilon\sim\mathcal{N}(0,1)$
7:         $\hat{x}_{0}=\frac{1}{{\sqrt{\bar{\alpha}_{t}}}}{(x_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}(x_{t},t))}$
8:         $x_{t-1}=\sqrt{\bar{\alpha}_{t-1}}\hat{x}_{0}+{\sqrt{1-\bar{\alpha}_{t-1}-\sigma_{t}^{2}}\epsilon_{\theta}(x_{t},t)+\sigma_{t}\epsilon}$
9:    end for
10:    $\hat{x}=\hat{x}+\hat{x}_{0}$
11:end for
12:$\hat{x}=\hat{x}/R_{t}$
13:Inverse Transform $\hat{x}$
14:return $\hat{x}$
@@ 3.3 Discussion
In this section, we first establish the connection between the diffusion model and the iterative denoising framework. Then we provide a rethinking of the diffusion model from a denoising perspective. Ultimately we point out the challenges in stimulating the diffusion model for image denoising. To begin with, the diffusion model and the iterative denoising framework share a similar structure. Formally they are all processes of iteratively generating high-quality images. In terms of details, both the sampling process eq. 3 and iteration eq. 5 can be decoupled into a prediction item $\hat{x}_{0}$ and an additional noise item. This additional noise is mainly a residual noise as $\epsilon_{\theta}(x_{t},t)$ in eq. 3 and $y-\hat{x}_{0}$ in eq. 5. It aims at introducing a weighted original noisy image. In summary, a single sampling process in diffusion models is equivalent to an iteration in the iterative denoising framework. The diffusion model and the iterative denoising framework both generate a roughly estimated image and introduce a weighted original noisy image $y$ to obtain a corrected version with lower noise. Currently, we have established that the diffusion model and the iterative denoising framework are similar in structure from a formal to a detailed level.
The functionality of a model is determined by its underlying structure. Thus, we believe the diffusion model can also perform denoising like the iterative denoising framework. With this in mind, we can examine the sampling process of the diffusion model from a denoising perspective. To explain this, suppose we start sampling from a given current image $x_{N}$. The image generated after the reverse process can be obtained by accumulating eq. 4:
@@
where $y=\frac{x_{N}}{\sqrt{\bar{\alpha}_{N}}}=x+\frac{\sqrt{1-\bar{\alpha}_{N}}}{\sqrt{\bar{\alpha}_{N}}}\epsilon$ and $\epsilon\sim\mathcal{N}(0,1)$. Here $x$ represents the desired clean image, and $y$ denotes its noisy observation.
The difference between a simple denoiser and eq. 6 lies in the cumulative terms. Here the cumulative terms represent the additional noise added and removed during the reverse process. During the reverse process, the network-estimated noise $\epsilon_{\theta}(x_{t},t)$ is subtracted to obtain the predicted clean image $\hat{x}_{0}$, and is added back to produce the next image $x_{t-1}$. Theoretically, the cumulative terms should be zero and not affect the level of denoising. In fact, the iterative denoising framework also has a similar cumulative term. In conclusion, a diffusion model can be considered as a denoiser, with its denoising ability at timestep $t$ given by ${\sqrt{1-\bar{\alpha}_{t}}}/{\sqrt{\bar{\alpha}_{t}}}$. Figure 1 illustrates the evolution of $x_{t}$, the predicted image $\hat{x}_{0}$, and the denoising level during the reverse process. In the early stages of the reverse process, the denoising level is high and the denoising effect is limited. So only low-frequency information can be determined. In the later stages, the remaining noise is minimal, thus high-frequency details can be refined. From a denoising perspective, the diffusion model can also be viewed as a coarse-to-fine iterative denoiser. These findings are exactly overlooked in previous research, and highlight the potential for leveraging a pre-trained unconditional diffusion model for image denoising.
Currently, we have identified the structural and functional similarities between the diffusion model and the iterative denoising framework. Now, we would like to discuss their differences. Firstly, the difference in structure between the diffusion model and the iterative denoising framework lies in the additional noise. The diffusion model, as shown in eq. 3, incorporates random Gaussian noise in its additional noise, whereas the Gaussian noise is absent in the iterative denoising framework. Secondly, the difference in functionality between the diffusion model and the iterative denoising framework pertains to their performances on distortion and perception. The iterative denoising framework primarily aims to reduce distortion, while the diffusion model excels in preserving high perception. The functionality of a model is determined by its underlying structure. Therefore, we believe that the key factors influencing distortion and perception are rooted in stochasticity.
The minor distinctions between the diffusion model and the iterative denoising framework do not undermine their structural similarity, given that both involve iterations comprising a prediction item and an additional noise item. The functionality of a model is determined by its underlying structure. Since their formal structures align, both the diffusion model and the iterative denoising framework can perform denoising in functionality. However, differences in stochasticity can influence denoising quality, impacting facets such as distortion and perception. We will delve deeper into our solutions for these challenges in Section 4.
@@
Finally, we have to address some critical issues when stimulating the diffusion model for image denoising tasks. One key issue is that input inconsistency. The pre-trained unconditional diffusion model is designed to accept pure Gaussian noise as input, whereas for denoising, the input is the noisy image. The existing restoration methods [48, 49] which mainly focus on other tasks, i.e.image super-resolution and image deblurring, begin their sampling process with Gaussian noise. As the sampling process advances from the initial state of pure noise to an intermediate state that generates an image $x_{N}$ with a noise level akin to that of the original input noisy image $y$, the content and details contained in $x_{N}$ do not align with those of the original input noisy image $y$. Consequently, existing methods not only exhibit poor performance but also demand a larger number of sampling times. In addition, to perform general denoising, we have to handle different noises. Thus, inspired by the iterative denoising framework, which starts from the noisy image, we propose an embedding method. We will find a way to embed the noisy image with different noise types into the sampling processes of the diffusion model. Another key issue is content inconsistency. The primary objective of the pre-trained unconditional diffusion model is the image synthesis task. While image synthesis just requires realistic images, image denoising requires high perception quality with minimal distortion. The excessive stochasticity of the diffusion model helps to generate realistic details which is beneficial to image synthesis. However, the inconsistent details between the generated images and the required denoised image will result in significant distortion. The existing restoration methods [48, 49] overlook this phenomenon and its underlying causes. We have identified the influential factors between distortion and perception and we will introduce an ensembling method to effectively adjust distortion and perception.

### v5

# Stimulating Diffusion Model for Image Denoising via Adaptive Embedding and Ensembling

@@ 1 Introduction
As smartphones have become ubiquitous, the pursuit of capturing high-quality images has become notably more demanding. Yet, when capturing images under challenging conditions, such as low-light environments, substantial information would be lost due to imaging noise. Consequently, the field of image denoising, with a focus on achieving low distortion and high perception, has been as a vital and thriving area of research.
Traditional methods [1, 2, 3] rely on image priors to guide the denoising process, but their effectiveness is limited under extreme conditions. With the development of deep learning, discriminative methods have become the mainstream method, which are usually trained by pixel-level losses [4, 5, 6, 7, 8]. Actually, the pixel-level losses tend to predict the median (or average) of all possible values rather than the realistic images [9, 10]. Thus, these discriminative methods often struggle with perceptual quality, particularly under extreme noise levels.
To improve perceptual quality, other image restoration tasks usually employ generative methods. Current state-of-the-art solutions usually rely on Generative Adversarial Networks (GANs) [11, 10, 12]. However, little work has been done to improve the perceptual quality of image denoising, even though generative approaches have been introduced to address real noise modeling [13, 14]. Moreover, GAN-based methods are widely blamed for artifacts and inconsistency, leading to significant distortion, especially for extreme degradation.
Figure 1: From the denoising perspective, the reverse process of a diffusion model can be viewed as a coarse-to-fine iterative denoising process. The noise level corresponds to the standard deviation of noise on an 8-bit image, where the maximum signal for a clean image is 255.
Recently, the diffusion model [15, 16, 17] has achieved state-of-the-art (SOTA) performance in various tasks, such as image super-resolution [18, 19], image inpainting [20, 21], image deblurring [22]. The methods [18, 20, 21, 22] that utilize diffusion models exhibit higher perceptual quality and fewer artifacts. Therefore, we believe that the diffusion model holds significant potential for image denoising. However, stimulating the diffusion model for image denoising still remains several critical problems. Specifically, the diffusion model is designed to receive standard Gaussian noise as input, which is different from the noisy image input required for image denoising. Furthermore, the content difference between the generated image and the desired denoised image introduces distortion. The input inconsistency problem and the content inconsistency problem are crucial to be tackled to stimulate the diffusion model for image denoising.
@@
In this paper, we propose a novel strategy to stimulate the Diffusion Model for Image Denoising (DMID), consisting of an adaptive embedding method and an adaptive ensembling method. Our insight comes from the denoising perspective towards the diffusion model. From the denoising perspective, the diffusion model and the iterative denoising methods share the same structure and similar functionalities. As illustrated in Figure 1, the diffusion model can be viewed as an iterative denoising method. This new perspective provides us with huge potential to employ a pre-trained unconditional diffusion model to tackle the aforementioned problems, which also eliminates the resource-intensive and time-consuming demands associated with training a diffusion model [23].
Firstly, the problem of input inconsistency is essentially an embedding problem. Here the noisy image and iterative denoising process correspond to the intermediate state and the reverse process subsequence of the pre-trained unconditional diffusion model, respectively. Therefore, we propose an adaptive embedding method that embeds the noisy image into an intermediate state of the pre-trained unconditional diffusion model. Our embedding method enables to perform Gaussian and real-world image denoising.
Secondly, the problem of content inconsistency fundamentally constitutes a stochasticity constraint problem. Since the distortion of the diffusion model mainly comes from the stochasticity inherent in the iterative process. To tackle this problem, we propose an adaptive ensembling method that can adjust distortion and perception without additional training. Our ensembling method enables to traverse through the perception-distortion curve [24] as shown in Figure 2, which achieves SOTA performance on both distortion-based and perception-based metrics.
In summary, our contributions are as follows:
@@
The organization of this paper is as follows. Firstly, we provide a brief introduction to the development of image denoising and the diffusion model in Section 2. Next, we offer a rethinking and understanding of the diffusion model from the denoising perspective in Section 3. Then, we propose our DMID strategy to tackle the existing problems in Section 4. After that, we demonstrate the performance and evaluate the effectiveness of our strategy in Section 5. Finally, we discuss the advantages and disadvantages in Section 6 and point out future directions in Section 7.
@@ 2.1 Image Denoising
Image denoising is an essential task in computer vision. It aims at restoring a clean image from its noisy counterpart. However, achieving high perception quality and low distortion in image denoising remains a challenging task. To tackle this problem, image denoising methods can be broadly divided into two categories: distortion-based methods and perception-based methods.
Distortion-based methods consist of traditional denoising methods and discriminative methods. Traditional image denoising methods focus on the usage of image priors such as sparsity [1, 25], low rank [2], self-similarity [26, 27], and smoothness [3, 28]. However, the ability of these priors in extreme conditions is ultimately limited, making it difficult to adequately denoise the corrupted images. With the development of deep learning, discriminative models [4, 5, 29, 30, 6, 7, 8, 31, 32], have become the mainstream methods. The main idea of discriminative methods is to use neural networks to learn mappings from noisy images to clean images through paired noisy-clean data or just noisy data [33, 34, 35, 36, 37]. $L_{1}$ loss and $L_{2}$ loss are most widely adopted during training. However, it is well known that these pixel-level losses tend to predict the median (or average) of all possible values rather than the realistic images [9, 10], resulting in over-smooth images with poor perceptual quality.
In addition, some distortion-based methods [38, 39, 40, 1] employ an iterative framework to optimize denoising performance. The iterative denoising methods usually predict a rough image and refine it through multiple iterations. The iterative denoising framework serves as a regularization and constraint. In addition, the iterative denoising framework decomposes image denoising into a series of sub-problems, making denoising easier. However, these methods still struggle with perceptual quality.
In contrast to distortion-based methods, perception-based methods typically exhibit superior perceptual quality but have significant distortion. Perception-based methods usually refer to generative methods. Current perception-based approaches with competitive performance for other image restoration usually rely on Generative Adversarial Networks (GAN) [11, 10, 12]. However, GAN-based methods often introduce artifacts and inconsistent details that are not present in the original clean image. The artifacts usually lead to significant distortion, especially for extreme degradation. As for image denoising, many methods employ GAN and flow [41] to solve real noise modeling problems [13, 42, 14, 43, 44, 45, 46] to get paired data easily. It is really a shame that little work has attempted to improve the perceptual quality of image denoising.
@@ 2.2 Diffusion Model
The diffusion model is a burgeoning likelihood-based generative model and has demonstrated remarkable success over other models in various tasks [23, 15, 47].
The diffusion model is originally proposed separately from diffusion-based [16] and score-matching-based [48] perspectives. Later Ho et al. [15] demonstrate the enormous potential of unconditional diffusion model for image synthesis. Recently Dhariwal et al.propose a kind of conditional diffusion model and bring diffusion models back to the public view. After that, conditional diffusion models adapt to various tasks and have achieved brilliant performances in a series of image restoration tasks, such as image super-resolution [18, 19], image inpainting [20, 21], image deblurring [22]. Some methods also exhibit the capability to address multiple tasks. For instance, the novel method [49] proposes a framework for training a diffusion model on a single image and is applicable in various tasks, including style transfer and harmonization. Another advanced method [21] exhibits proficiency in tasks such as super-resolution, deblurring, and inpainting.
Recently, some methods [50, 51, 21] have exclusively employed pre-trained unconditional diffusion models, originally trained for image synthesis, to address linear image restoration problems. These methods typically assume that the observed image $y$ is degraded as $y=Hx+n$, where $H$ represents the degradation matrix, $n$ denotes noise, and $x$ signifies the desired clean image. The existing methods usually replace some content with the original degraded image and preserve the others during each sampling step of the diffusion model. These methods achieve this by decomposing the degradation matrix $H$ to identify the boundary of the reserved information area. The popular techniques are Singular Value Decomposition and Range-Null Space Decomposition.
@@
More recently, a method [52] attempts to customize different diffusion models for each noise type. However, to customize distinct diffusion models, the method has to simplify noise types to standard Gaussian or standard Poisson noise, making it less effective for real-world image denoising.
@@ 3.1 Diffusion Model Pipeline
In this section, we give a rough introduction to the diffusion model.
Diffusion models are composed of a T-timestep forward process and a T-timestep backward process. The forward process gradually adds Gaussian noise to the initial data $x_{0}$ and converges to the standard Gaussian distribution. The reverse process starts with standard Gaussian noise $x_{T}\sim\mathcal{N}(0,1)$ and eventually obtains the expected high-quality image $x_{0}$.
The forward process is changed from the previous state $x_{t-1}$ follows the Markov chain to generate current state $x_{t}$:
@@
where $\alpha_{t}=1-\beta_{t}$, $\bar{\alpha_{t}}=\prod_{i=1}^{t}(1-\beta_{i})$.
The reverse process generates an image by a series of sampling processes, and we refer to the number of sampling times as $S_{t}$. For instance, from the initial state $x_{1000}$ to get the intermediate state $x_{500}$ and from the intermediate state $x_{500}$ to obtain the final state $x_{0}$, the sampling times are $S_{t}=2$. The sampling process from $x_{t}$ to $x_{t-1}$ is expressed as:
@@
where $\mu_{\theta}(x_{t},t)$ is modeled by the neural network. The reverse process gradually transforms the Gaussian distribution $x_{T}$ into the expected data distribution $x_{0}$.
Different sampling strategies [15, 54] of Eq. 3 can be summarized as a predicted item and an additional noise item:
@@
3.2 Iterative Denoising Framework
In this section, we point out the iterative framework for image denoising.
Some distortion-based methods [38, 39, 40, 1] employ an iterative framework to optimize denoising performance. These methods usually predict a rough image and refine it through many iterations. Such an iterative framework breaks down the problem into a series of sub-problems, leading to a coarse-to-fine denoising process. The iterative denoising framework can be developed as Algorithm 1.
Specifically, the inference of this framework includes several iterations. In each iteration, a rough denoising result $\hat{x}_{0}$ is firstly estimated from the current image $x_{t}$. After that, a weighted original noisy image $y$ is introduced to obtain a corrected version $x_{t-1}$ with lower noise than the previous image $x_{t}$, as:
$x_{t-1}=\hat{x}_{0}+\gamma_{t}(y-\hat{x}_{0})\ .$ (5)
@@ 3.2 Iterative Denoising Framework
2:$k=\text{floor}({N/S_{t}})$
3:for $i=1,\dotsc,R_{t}$ do
4:    for $t$ in reversed(range$(0,N,k)$do
5:         $\epsilon\sim\mathcal{N}(0,1)$
6:         $\hat{x}_{0}=\frac{1}{{\sqrt{\bar{\alpha}_{t}}}}{(x_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}(x_{t},t))}$
7:         $x_{t-k}=\sqrt{\bar{\alpha}_{t-k}}\hat{x}_{0}+{\sqrt{1-\bar{\alpha}_{t-k}-\sigma_{t}^{2}}\epsilon_{\theta}(x_{t},t)+\sigma_{t}\epsilon}$
8:    end for
9:    $c_{i}=\hat{x}_{0}$
10:end for
11:$\hat{x}=\sum_{i=1}^{R_{t}}c_{i}p(y|c_{i})/\sum_{i=1}^{R_{t}}p(y|c_{i})$
12:Inverse Transform $\hat{x}$
13:return $\hat{x}$
@@ 3.3 Discussion
Denoising Perspective Understanding. Form follows function [55], which is a well-known law, indicating that similar forms can support similar functions. The high formal similarity between the diffusion model and the iterative denoising framework inspires us to explore their relevance.
The diffusion model and the iterative denoising framework are similar from structures to details. In terms of structures, both techniques build a coarse-to-fine iterative process to generate high-quality images. Each step produces a better iteration result by the weighted fusion of coarse denoised image $\hat{x}_{0}$ and noisy image $y$. In terms of details, both the sampling process Eq. 4 and iteration Eq. 5 can be decoupled into a prediction item $\hat{x}_{0}$ and an additional noise item. This additional noise is mainly a residual noise as $\epsilon_{\theta}(x_{t},t)$ in Eq. 4 and $y-\hat{x}_{0}$ in Eq. 5. Therefore, we declare that the diffusion model and the iterative denoising framework are indeed similar in form, where one sampling step of the former corresponds to one iteration of the latter.
Based on the above observations, we can rewrite the sampling process of the diffusion model from a denoising perspective. Suppose that we start sampling from the intermediate state $x_{N}$ at timestep $N$. The generated image $x_{0}$ after $N$ steps of reverse process can be obtained by accumulating Eq. 4:
@@
where $y=\frac{x_{N}}{\sqrt{\bar{\alpha}_{N}}}=x+\frac{\sqrt{1-\bar{\alpha}_{N}}}{\sqrt{\bar{\alpha}_{N}}}\epsilon$ and $\epsilon\sim\mathcal{N}(0,1)$. Here $x$ represents the desired clean image, and $y$ denotes its noisy observation. The derivation of Eq. 6 is given in the supplementary material.
The difference between a simple denoiser and Eq. 6 lies in the cumulative terms. Here the cumulative terms represent the additional noise added and removed during the reverse process. During the reverse process, the network-estimated noise $\epsilon_{\theta}(x_{t},t)$ is subtracted to obtain the predicted clean image $\hat{x}_{0}$, and is added back to produce the next image $x_{t-1}$. Theoretically, the cumulative terms should be zero and not affect the level of denoising. In fact, the iterative denoising framework also has a similar cumulative term.
Overall, we introduce a novel denoising perspective for understanding the diffusion model. The diffusion model can be classified as a variant of the iterative denoising framework. The neural network of the diffusion model can be regarded as a denoiser, where timestep $t$ as a variable corresponds to the preset noise level ${\sqrt{1-\bar{\alpha}_{t}}}/{\sqrt{\bar{\alpha}_{t}}}$.
Figure 1 illustrates the evolution of the intermediate state $x_{t}$, the estimated image $\hat{x}_{0}$, and the corresponding noise level during the reverse process. In the early stages of the reverse process, the noise level is high and the denoising effect is limited. So only low-frequency information can be determined. In the later stages of the reverse process, the remaining noise is minimal, and then high-frequency details are generated with iterations.
Denoising perspective understanding will intuitively and clearly guide us to stimulate the diffusion model for image denoising. To the best of our knowledge, this is the first time that the diffusion model is comprehended in such a concise denoising perspective.
@@
Denoising Perspective Rethinking. The denoising perspective of the diffusion model is the starting point for us to rethink and stimulate the diffusion model for image denoising. The input of diffusion models is standard Gaussian noise while the input of image denoising is a noisy image, which constitutes input inconsistency. The content difference between the generated image and the desired denoised image introduces distortion, which constitutes content inconsistency. Guided by the denoising perspective, the approach to addressing input inconsistency and content inconsistency naturally emerges.
For input inconsistency, the solution lies in the comparison between the input of the diffusion model and the iterative denoising framework. From a denoising perspective, the diffusion model is a conditional AWGN denoiser, iteratively denoising according to a schedule where the expected noise level of the intermediate state $x_{t}$ gradually decreases. Therefore, by proposing a method to embed the input into the corresponding intermediate state of the pre-trained diffusion model based on the noise level, we naturally address input inconsistency.
For content inconsistency, the solution lies in the comparison between the process of the diffusion model and the iterative denoising framework. The diffusion model excels in preserving high perceptual quality, while the classical iterative denoising framework primarily aims to reduce distortion. Based on our observations, the difference is mainly attributed to the introduction of random noise $\epsilon$ besides the original noisy image $y$ during the sampling process of the diffusion model. The excessive stochasticity of random noise $\epsilon$ allows the diffusion model to potentially generate content beyond the original clean image $x_{0}$. From the denoising perspective, this phenomenon corresponds to the local content flickering caused by the stochasticity of i.i.d. noise in low-light video denoising [56]. Taking cues from the late fusion in video processing, we introduce an ensembling method to effectively adjust distortion and perception in the diffusion model, thereby addressing content inconsistency.
It is worth highlighting that our method is insightful compared to existing diffusion-based image denoising methods [50, 51]. Existing methods are not specifically designed for image denoising, thus they typically start from pure Gaussian noise and recommend extremely high sampling times for one inference , which is time-consuming and inefficient. The denoising perspective helps us rethink the diffusion model and addresses the problems of the input inconsistency and the content inconsistency in a concise and efficient manner.